# AttributionScanner: A Visual Analytics System for Model Validation with Metadata-Free Slice Finding

### v3

@@ I Introduction
However, models often perform inconsistently across different subsets of data, making it challenging to validate a model comprehensively to uncover hidden issues [4, 5, 6, 7, 8]. Thus, the crucial demands of identifying and characterizing such problematic data subsets have fueled data slice finding techniques [4, 5, 6, 7, 8, 9, 10]. The term “data slice” refers to data subgroups with coherent features, often defined by metadata, model representations, or other descriptive information. By detecting, understanding, and resolving issues in data slices, ML experts can improve model performance and ensure safer and more reliable deployments in high-stakes domains [11, 12, 13, 14]. To enhance the trustworthiness, eXplainable Artificial Intelligence (XAI) techniques [15] are also utilized to reveal potential factors adversely impacting the model and apply targeted model regularization [16].
@@
Challenges of Data Slice Finding. State-of-the-art methods for data slice finding demand extensive metadata associated with each single instance, which are often obtained from manual human annotations[17, 18, 19] or pre-trained vision-language models[14]. However, it is expensive to collect manually annotated metadata, and vision-language models can produce erroneous results, especially in domain-specific scenarios. When neither metadata nor suitable pre-trained models are available, obtaining data slices becomes challenging: requires manual examination of numerous instances or computationally expensive model training[20, 21].
@@
Our approach. We present AttributionScanner, a Visual Analytics (VA) system designed for efficient vision model validation with slice finding, eliminating the costly requirements of metadata or other descriptive information. Our innovative workflow employs attribution-based explanation techniques to create interpretable feature representations for slice finding, enabling discriminative slice construction and intuitive data summary to streamline user exploration. Additionally, the methods provided by AttributionScanner are applicable to various computer vision models, collaboratively forming a human-in-the-loop approach for model validation and enhancement. Our main contributions include:
@@
This paper is structured as follows: in Sec. II, we review related work on data slice finding and XAI that is relevant to our work Sec. III outlines the methodology and system design of AttributionScanner. In Sec. IV, we illustrate two use cases showcasing how AttributionScanner can be utilized for vision model validation. An evaluation of our system and experts’ feedback are discussed in Sec. V. Lastly, we present our future work and conclusions in Sec. VI.
@@ II-A Data Slice Finding
Considering the fact the ML models often perform inconsistently across different data subsets, data slice finding technique is used to identify data subsets with coherent patterns, which can be defined by similar metadata, model representations, data statistics, or other descriptive information [26, 17, 9, 18]. By analyzing problematic data slices, ML experts can design targeted solutions to address model issues in edge cases, improving the model robustness and reliability more efficiently [4]. Existing approaches often necessitate expensive metadata to obtain meaningful subgroups, which hinders the applicability of data slice finding. To address this, a prevalent approach leverages feature vectors in the latent space and then clusters them to produce data slices [20, 21, 27, 28]. To provide data slice explanation, Domino [14] deploys a pre-trained vision-language model (VLM) to generate textual descriptions of slices, whose reliance on VLM can backfire if domain mismatch is significant, yielding less-meaningful slices. To fill the gap, we introduce a system providing explainable slices without metadata or pre-trained VLMs.
@@ II-B Visual Analytics for Model Explanation and Validation
Visual analytics have also proven effective for model explanation [29, 30, 31, 32, 33, 34, 35] and validation [36, 37, 38, 39, 40]. Using pairwise feature combinations, [41] produces data slices and assesses model performance with data cube analysis method. FairVis [42] enables users to manually slice and inspect data slices based on domain knowledge to identify potential model biases. Errudite [43] develops a domain-specific language for data slicing of textual documents. Although manually slicing the data is useful to a certain extent, it is not scalable because of its expansiveness. To address this issue, SliceTeller [10] combined an automatic slice finding algorithm with a visual analytics tool to allow the analysis and iterative refinement of models. However, it still required structured metadata to produce meaningful data slices. Our approach is for explainable model validation to eliminate the expensive metadata while involving effective human participation.
Figure 1: AttributionScanner workflow involves three phases: Explainable Data Slice Finding, Slice Summarization and Annotation, and Slice Error Mitigation. The first phase: GradCAM is used to assist the generation of feature vectors and then obtain data slices. The second phase: users can identify and annotate slice error types such as core/spurious correlations and noisy labels with the help of Attribution Mosaic and Spuriousness propagation. The third phase: the annotation and user-verified Spuriousness are used on the ML side to mitigate slice errors, including (1) Data re-labeling if considered necessary, and (2) noise corruption and Core Risk Minimization (CoRM) applied if spurious correlations are detected.
@@ II-C XAI Techniques for Neural Networks
Existing XAI techniques often use visualizations to support neural network interpretation, providing attribution heatmaps [22, 23, 24, 44] or visualizing models’ learned patterns [45, 46, 47]. Specifically, attribution-based explanations typically utilize information such as model gradients [23, 24], decision boundaries [22], or causal relationships [25, 44] to explain the model’s specific decision. However, despite the capability of offering intuitive and straightforward results, these methods are instance-based, explaining single data instances and missing the big picture of the common model behavior. On the other hand, methods such as Feature Visualization [45, 46] and Feature Inversion [47] create visualizations that maximally activate specific model neurons or layers, or visualizations that lead to feature vectors mostly similar to a target, which can be an aggregation of a group of feature vectors. In our approach, we leverage the benefit of Feature Inversion to highlight the shared patterns of a group of instances for visual summaries of our data slices, and also deploy GradCAM, a widely adopted attribution-based explanation method for explaining single instances.
@@ III AttributionScanner
In this section, we introduce AttributionScanner, a novel approach for interpretable ML model validation with metadata-free, data slice-level analysis. To foster a thorough understanding of our system, we first outline the requirements that shaped our design choices. Then, we delve into the visual components of AttributionScanner that enable interpretable model validation. We then detail our slice computation and summarization method, which allow us to generate data slices with consistent model attribution without additional metadata. Finally, we show how we leverage the insights obtained from our system to enhance the model.
@@ III-B System Workflow
A workflow of our approach is presented in Fig. 1, where we foster explainable slice finding without metadata ( R1, R2) by leveraging XAI technique GradCAM [24] and Feature Inversion [46]. The first “explainable data slice finding” phase interpolates model attributions from GradCAM into the latent space to craft attribution-weighted feature vectors, which enable us to form an attribution representation space with locally consistent model attributes ( R1). Then we conduct clustering over this space to generate data slices. Although the produced slices maintain consistent attributions, they still demand significant human effort to identify and understand their shared patterns. To streamline slice exploration and issue detection, the “slice summarization and annotation” phase introduces Attribution Mosaic, which visually elucidates each slice’s dominant attributions ( R3), synchronized with instance-level heatmaps for insight verification ( R2). Upon user annotation of an issue like “spurious correlations”, our Spuriousness Propagation method automatically estimates a spuriousness score for each slice to help users uncover other hidden model biases ( R4). Afterward, our workflow enables the mitigation of detected issues in data or model, corresponding to the “slice error mitigation” phase ( R4).
The rest of this section delineates our interface for data slicing-based model validation, explains our presented data slicing and summarization approach, discusses design considerations, and demonstrates how the model can be fine-tuned to rectify detected slice issues.
Figure 2: AttributionScanner applied to the model validation of a hair color classifier trained on the CelebA dataset. A System Menu, enabling the selection of dataset, model, and visualization options (Layout, Confusion Matrix, Scatter Plot/Attribution Mosaic, and Contour Visibility). B Slice Table, showing slice metrics. C Attribution Mosaic, showing a visual overview of all data slices, which can also be displayed as a confusion matrix view ( E ). D Slice Detail View, showing individual images or Attribution Heatmaps belonging to a selected data slice.
@@ III-C System Interface
Figure 3: Attribution-weighted feature vector generation. An image is forwarded through CNN, where the corresponding feature vector $F$ and weight matrix $W$ can be extracted. Through linear interpolation of $W$ to the image space, we obtain the attribution heatmap $A$ to calculate the attribution-weighted feature vector $F_{W}$.
@@ III-D Explainable Data Slice Finding
The process unfolds in three phases: initially, we create attribution-weighted feature vectors utilizing attribution masks in the latent space. Subsequently, we derive an attribution representation space and compute data slices. Lastly, we employ the Feature Inversion [47] to distill shared visual patterns per slice and visualize these patterns alongside model performance indicators in Attribution Mosaic.
(1) Attribution-Weighted Feature Vector Generation
@@
Figure 4: Comparison of the representation spaces for the hair color classification model. (a) Feature representation space computed on original feature vectors. (b) Attribution representation space computed on attribution-weighted feature vectors.
(2) Data Slice Identification
Representation Space Construction. To derive meaningful data slices ( R1, R2), we need to construct a representation space with neighbor consistency in terms of model attributes and data features. We first conduct quantitative and qualitative experiments to validate whether attribution-weighted feature vectors help fulfill our requirement. Specifically, we compare the feature representation space (derived from the original feature vectors) and the attribution representation space (derived from the attribution-weighted feature vectors) As shown in Fig. 4, following the established practices[14, 4, 5] and considering the time efficiency [58, 59], we utilize dimensionality reduction by UMAP [58] to produce both spaces. At each space, three points are randomly sampled and their nearest neighbors are obtained. We qualitatively examine the consistency of their corresponding attribution heatmaps. As showcased in Fig. 4, neighboring instances’ model attributions on the original space are not similar, while our constructed attribution representation space has greater local consistency concerning model attributions.
@@
Data Slice Identification. On top of the attribution representation space, we apply K-Means clustering to obtain data slices. Similarly to DOMINO [14], we apply over-clustering by incrementing the number of clusters and monitoring the model attribution consistency within each slice, until a coherent grouping of samples is attained. Specifically, the attribution consistency of a data slice is measured as the average cosine similarity of each instance’s attribution-weighted feature vector with the slice centroid. We set a threshold at $0.8$ to determine slice coherence in our case studies. Note that we allow ML experts to adjust this threshold based on the specific requirements of their datasets and tasks. This helps prevent the omission or overlooking of rare slices and provides a comprehensive slice identification.
@@ III-E Slice Summarization and Annotation
(1) Attribution Mosaic
After obtaining slices with consistent model attributions, we design Attribution Mosaic, a novel visualization technique, to visually summarize data slices in a mosaic depiction( R3). We present the pipeline in Fig. 5. First, we compute the convex hull of each slice based on the projected points in the 2D space. We then distill each slice’s predominant visual patterns via Feature Inversion, integrating the shape constraint of the convex hull. The final landscape featuring slices’ attribution summaries is presented as Attribution Mosaic.
Mosaic Boundary Generation. Given that the convolutional modules take rectangular-shaped input, conventional Feature Inversion produces rectangular visualizations, merely obtaining and displaying its results over the projected cluster centroids leads to overlapped visualizations, hindering effective user exploration. Since K-Means is a centroid-based algorithm that partitions the space into strict non-overlapping Voronoi cells, this guarantees that the convex hulls [60] of our data slices have no overlap.
Figure 5: Attribution Mosaic generation. After acquiring attribution-weighted feature vectors ($F_{W}$), we compute data slices within the attribution representation space. Then Feature Inversion is conducted according to $F_{W}$ and the mosaic boundary of each slice to visualize their common patterns, forming Attribution Mosaic.
Therefore, we compute the convex hull of each slice on the attribution representation space, which provides a mosaic layout for our slice visualization.
@@
(2) Slice Annotation and Spuriousness Propagation
We introduce a metric termed Spuriousness probabilities to streamline the detection of spuriousness, ranging from 00 to $1$, indicative of the likelihood of spurious correlation existence. We employ the Label Spreading algorithm [64] to propagate user-annotated spurious/core information to other slices. Specifically, a graph is constructed where nodes represent data points and edge weights represent similarities between points. In our case, we fit the graph to our 2D attribution representation space.
@@
According to the graph, the Label Spreading algorithm updates the labels of unannotated slices iteratively to ensure local and global label consistency—similar slices should have similar spuriousness, and the propagated spuriousness scores should respect the annotated slice spurious/core labels. The minimal input requirement from users is one slice annotation of “core” or “spurious” (Spuriousness=00 or $1$), and then AttributionScanner will propagate this information to create Spuriousness probabilities for the rest of slices.
The Spuriousness probabilities obtained through the Attribution Mosaic offer two benefits. First, they streamline slice exploration by highlighting potential spurious correlations, supporting identifying and assessing problematic slices. Second, after users’ verification, these probabilities inform the model-side strategies for issue mitigation (Sec. III-G) ( R4).
@@ IV-A Hair Color Classifier Validation - Finding Edge Cases
$\bullet$ Does the model behave correctly on well-performing slices? One basic expected behavior for a hair color classification model is to catch hair features. At first glance of the Attribution Mosaic (Fig. 2  C), ML experts notice slice_22 and slice_5 on the left, which lie separately with others on the Attribution Mosaic.
@@
Their mosaic visualizations indicate gray hair patterns, which means the model uses the correct features, i.e., core features. By verifying the corresponding attribution heatmaps (Fig. 2  D), they confirm the correctness of this insight and annotate both slices as core feature with description Core: gray hair. Upon experts saving the annotation, AttributionScanner automatically propagates the annotation and provides a Spuriousness probability of each slice. With this guidance, the experts observe many slices lying on the right of Attribution Mosaic have higher Spuriousness probabilities (Fig. 2  C  E). Their mosaic visualizations do not show any hair patterns, leading to valid doubts that the model does not behave correctly on such slices. Besides, they also notice that the model has correct predictions on these slices (with $100\%$ prediction accuracy). This makes them worry that the model is largely biased by spurious features. Through investigation, they find the model mistakenly utilizes mouth and eyes to predict hair color for those top-performed slices, such as slice_25, slice_14, and slice_2 (Fig. 9).

### v4

@@ I Introduction
However, models often perform inconsistently across different subsets of data, making it challenging to validate a model comprehensively to uncover hidden issues [4, 5, 6, 7, 8]. Thus, the crucial demands of identifying and characterizing such problematic data subsets have fueled data slice finding techniques [4, 5, 6, 8, 9, 10]. The term “data slice” refers to data subgroups with coherent features, often defined by metadata, model representations, or other descriptive information. By detecting, understanding, and resolving issues in data slices, ML experts can improve model performance and ensure safer and more reliable deployments in high-stakes domains [11, 12, 13, 14]. To enhance the trustworthiness, eXplainable Artificial Intelligence (XAI) techniques [15] are also utilized to reveal potential factors adversely impacting the model and apply targeted model regularization [16].
@@
Challenges of Data Slice Finding. State-of-the-art methods for data slice finding demand extensive metadata associated with each single instance, which are often obtained from manual annotations [17, 18, 19] or generated by pre-trained vision-language models [14]. However, it is expensive to collect manually annotated metadata, and vision-language models can produce erroneous results [20], especially in domain-specific scenarios. When neither metadata nor suitable pre-trained models are available, obtaining data slices becomes challenging: requires manual examination of numerous instances or computationally expensive model training [21, 22].
@@
Our approach. We present AttributionScanner, a Visual Analytics (VA) system designed for efficient vision model validation with slice finding, eliminating the costly requirements of metadata or other descriptive information. Our innovative workflow employs attribution-based explanation techniques to create interpretable feature representations for slice finding, enabling discriminative slice construction and summary to streamline user exploration. Additionally, the methods provided by AttributionScanner are applicable to various computer vision models, collaboratively forming a human-in-the-loop approach for model validation and enhancement. Our main contributions include:
@@ II-A Data Slice Finding
Data slice finding is a critical technique for identifying coherent subsets of data where ML models may perform inconsistently. These subsets, or slices, are often defined by similar metadata, model representations, data statistics, or other descriptive information [27, 17, 9, 18]. By analyzing problematic data slices, ML experts can design targeted solutions to address model issues in edge cases, improving the model robustness and reliability [4, 7]. However, existing approaches often rely on expensive metadata to obtain meaningful subgroups, limiting their applicability. To address this, recent work has focused on leveraging feature vectors in the latent space and clustering them to produce data slices [21, 22, 28, 29]. For example, Domino [14] uses a pre-trained vision-language model (VLM) to generate textual descriptions of slices. However, this approach can backfire if there is a significant domain mismatch, yielding less meaningful slices. Our work fills this gap by introducing a system that provides explainable slices without requiring metadata or pre-trained VLMs, making it more broadly applicable and reliable.
@@ II-B Visual Analytics for Model Explanation and Validation
Visual analytics has emerged as a powerful tool for both model explanation [30, 31, 32, 33, 34, 35, 36] and validation [37, 38, 39, 40, 41, 42]. For instance, Kahng et al. [43] use pairwise feature combinations to produce data slices and assess model performance. FairVis [44] enables domain users to manually slice data to identify model biases. Similarly, Errudite [45] develops a domain-specific language for slicing textual documents. While manual slicing is useful, it is not scalable due to its labor-intensive nature. To address this, SliceTeller [10] combines an automatic slice-finding algorithm with a visual analytics tool, allowing for iterative refinement of models. However, it still relies on structured metadata to produce meaningful slices. Our approach, AttributionScanner, eliminates the need for expensive metadata while incorporating effective human participation, making it a scalable and accessible solution for explainable model validation.
Figure 1: AttributionScanner workflow involves three phases: Explainable Data Slice Finding, Slice Summarization and Annotation, and Slice Error Mitigation. The first phase: GradCAM is used to assist the generation of feature vectors and then obtain data slices. The second phase: users can identify and annotate slice error types such as core/spurious correlations and noisy labels with the help of Attribution Mosaic and Spuriousness propagation. The third phase: the annotation and user-verified Spuriousness are used on the ML side to mitigate slice errors.
@@ II-C XAI Techniques for Neural Networks
Explainable AI (XAI) techniques have become essential for interpreting neural networks, often using visualizations to provide insights into model behavior. Attribution-based methods [24, 25, 23, 46, 26], such as GradCAM [25], generate heatmaps or decision boundary explanations for individual instances. While these methods offer intuitive and straightforward results, they are instance-based and fail to capture broader patterns in model behavior. Another line of work utilizes optimization to produce synthetic images for model explanation, such as Feature Visualization [47, 48] and Feature Inversion [49], which create visualizations that maximally activate specific neurons/layers, or reconstruct images from feature vectors. These methods provide insights into the learned patterns of a model but are often limited to single instances or randomly aggregated groups. In AttributionScanner, we bridge this gap by leveraging Feature Inversion to highlight shared patterns across data slices, providing visual summaries of model behavior at the slice level. Additionally, we integrate GradCAM to explain individual instances, offering a comprehensive view of model behavior at both the micro and macro levels.
@@ III-B System Workflow
A workflow of our approach is presented in Fig. 1, where we foster explainable slice finding without metadata ( R1, R2) by leveraging GradCAM [25] and Feature Inversion [48]. The first “explainable data slice finding” phase interpolates model attributions into the latent space to craft attribution-weighted feature vectors, forming an attribution representation space with locally consistent model attributes ( R1). Then we conduct clustering over this space to generate data slices. Although the produced slices maintain consistent attributions, they still demand significant human effort to identify and understand their shared patterns. To streamline slice exploration and issue detection, the “slice summarization and annotation” phase introduces Attribution Mosaic, which visually elucidates each slice’s dominant attributions ( R3), synchronized with instance-level heatmaps for insight verification ( R2). Upon user annotation of an issue like “spurious correlations”, our Spuriousness Propagation method automatically estimates a spuriousness score for each slice to help users uncover other hidden model biases ( R4). Afterward, our workflow enables the mitigation of detected issues in data or model, corresponding to the “slice error mitigation” phase ( R4).
Figure 2: AttributionScanner applied to the model validation of a hair color classifier trained on the CelebA dataset. (A) System Menu, enabling the selection of dataset, model, and visualization options (Layout, Confusion Matrix, Scatter Plot/Attribution Mosaic, and Contour Visibility). (B) Slice Table, showing slice metrics. (C) Attribution Mosaic, showing a visual overview of all data slices, which can also be displayed as a confusion matrix view (E). (D) Slice Detail View, showing individual images or Attribution Heatmaps belonging to a selected data slice.
@@ III-C System Interface
Figure 3: Attribution-weighted feature vector generation. An image is forwarded through CNN, where the corresponding feature vector $F$ and weight matrix $W$ can be extracted to calculate the attribution-weighted feature vector $F_{W}$.
@@ III-D Explainable Data Slice Finding
III-D1 Attribution-Weighted Feature Vector Generation
@@
Figure 4: Comparison of representation spaces. (a) Feature representation space computed on original feature vectors. (b) Attribution representation space computed on attribution-weighted feature vectors.
III-D2 Data Slice Identification
Representation Space Construction. To derive meaningful data slices ( R1, R2), we need to construct a representation space with neighbor consistency in terms of model attributes and data features. We first conduct quantitative and qualitative experiments to validate whether attribution-weighted feature vectors help fulfill this requirement—by comparing the feature representation space (derived from the original feature vectors) and the attribution representation space (derived from the attribution-weighted feature vectors). Following the established practices [14, 4, 5] and considering the time efficiency [60, 61], we utilize dimensionality reduction by UMAP [60] to produce both spaces. At each space, three points are randomly sampled and their nearest neighbors are obtained. We qualitatively examine the consistency of their corresponding attribution heatmaps as shown in Fig. 4. Neighboring instances’ model attributions on the original space are not similar, while our constructed space has greater local consistency of model attributions.
@@ III-D2 Data Slice Identification
Slice Identification. On top of the attribution representation space, we apply K-Means clustering to obtain data slices. Similarly to Domino [14], we apply over-clustering by incrementing the number of clusters and monitoring the model attribution consistency within each slice, until a coherent grouping of samples is attained. Specifically, the attribution consistency of a data slice is measured as the average cosine similarity of each instance’s attribution-weighted feature vector with the slice centroid. We set a threshold at $0.8$ to determine slice coherence in our case studies.
Figure 5: Attribution Mosaic generation. We compute data slices after acquiring attribution-weighted feature vectors ($F_{W}$). Then Feature Inversion is conducted according to $F_{W}$ and the mosaic boundary of each slice to visualize their common patterns, forming Attribution Mosaic.
Note that we allow ML experts to adjust this threshold based on the specific requirements of their datasets and tasks. For rare slices with a small number of data samples featuring abnormal model attributions, this iterative process, along with our slice mosaic design detailed in Sec. III-E, ensures such slices are not overlooked. By combining automated clustering with user-driven refinement, our approach provides a comprehensive and efficient solution for identifying both common and rare slices.
@@ III-E Slice Summarization and Annotation
III-E1 Attribution Mosaic
After obtaining slices with internally consistent model attributions and covering the entire dataset, we design Attribution Mosaic, a novel visualization technique, to visually summarize data slices in a mosaic depiction( R3). As shown in Fig. 5, we first compute the convex hull of each slice based on the projected points in 2D space; then distill each slice’s predominant visual patterns via Feature Inversion, integrating the shape constraint of the convex hull. The final landscape featuring slices’ attribution summaries is presented as Attribution Mosaic, allowing users to easily identify slices with rare, or uncommon model attributions through visual inspection and comparison.
Mosaic Boundary Generation. Given that the convolutional modules take rectangular-shaped input, conventional Feature Inversion produces rectangular visualizations, merely obtaining and displaying its results over the projected cluster centroids leads to overlapped visualizations, hindering effective user exploration. Since K-Means is a centroid-based algorithm that partitions the space into strict non-overlapping Voronoi cells, this guarantees that the convex hulls [62] of our data slices have no overlap. Therefore, we compute the convex hull of each slice on the attribution representation space, which provides a mosaic layout for our slice visualization.
@@ III-E1 Attribution Mosaic
III-E2 Slice Annotation and Spuriousness Propagation
We introduce a metric termed Spuriousness probabilities to streamline the detection of spuriousness, ranging from 00 to $1$, indicative of the likelihood of spurious correlation existence. We employ the Label Spreading algorithm [66] to propagate user-annotated spurious/core information to other slices. Specifically, a graph is constructed where nodes represent data points, and edge weights encode similarities between points. In our case, we fit the graph to our 2D attribution representation space. Based on this graph structure, the Label Spreading algorithm iteratively updates the labels of unannotated slices, ensuring local and global label consistency—similar slices should exhibit similar spuriousness scores, and the propagated scores should respect the annotated slice spurious/core labels. While only one annotation of “core” or “spurious” is required to initialize the propagation process, the spuriousness scores are dynamically updated as users provide additional annotations. The iteratively refined scores are visualized in the interface to guide further annotation.
The Spuriousness probabilities obtained through the Attribution Mosaic offer two benefits. First, they streamline slice exploration by highlighting potential spurious correlations, supporting identifying and assessing problematic slices. Second, after users’ verification, these probabilities inform the model-side strategies for issue mitigation (Sec. III-G) ( R4). Since propagated scores are visualized alongside the Attribution Mosaic, users can easily validate and override incorrect propagations, ensuring human decisions take precedence over automated computations.
@@ IV-A Hair Color Classifier Validation - Finding Edge Cases
$\bullet$ Does the model behave correctly on well-performing slices? One basic expected behavior for a hair color classification model is to catch hair features. At first glance of the Attribution Mosaic (Fig. 2  C), ML experts notice slice_22 and slice_5 on the left, which lie separately with others on the Attribution Mosaic. Their mosaic visualizations indicate gray hair patterns, which means the model uses the correct features, i.e., core features. By verifying the corresponding attribution heatmaps (Fig. 2  D), they confirm the correctness of this insight and annotate both slices as core feature with description Core: gray hair. Upon experts saving the annotation, AttributionScanner automatically propagates the annotation and provides a Spuriousness probability of each slice. With this guidance, the experts observe many slices lying on the right of Attribution Mosaic have higher Spuriousness probabilities (Fig. 2  C  E). Their mosaic visualizations do not show any hair patterns, leading to valid doubts that the model does not behave correctly on such slices. Besides, they also notice that the model has correct predictions on these slices (with $100\%$ prediction accuracy). This makes them worry that the model is largely biased by spurious features. Through investigation, they find the model mistakenly utilizes mouth and eyes to predict hair color for those top-performed slices, such as slice_25, slice_14, and slice_2 (Fig. 9).